# XPUTimer: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale

### v1

@@ Abstract
The rapid proliferation of large language models has driven the need for efficient GPU training clusters. However, ensuring high-performance training in these clusters is challenging due to the complexity of software-hardware interactions and the frequent occurrence of training anomalies. Since existing diagnostic tools are narrowly tailored to specific issues, there are gaps in their ability to address anomalies spanning the entire training stack. In response, we introduce XPUTimer, a real-time diagnostic framework designed for distributed LLM training at scale. XPUTimer first integrates a lightweight tracing daemon to monitor key code segments with minimal overhead. Additionally, it features a diagnostic engine that employs novel intra-kernel tracing and holistic aggregated metrics to efficiently identify and resolve anomalies. Deployment of XPUTimer across 6,000 GPUs over eight months demonstrated significant improvements across the training stack, validating its effectiveness in real-world scenarios.
@@ 1 Introduction
The advent of large language models (LLMs) has revolutionized the deep learning training community, driving substantial advancements in artificial intelligence-generated content (AIGC). Recognizing their transformative potential to enhance user experiences, leading corporations are proactively leveraging LLMs to enhance a wide array of user-oriented services [1, 2, 3]. To meet the computational demands of LLM training, they construct large-scale training clusters comprising the latest GPUs interconnected via high-bandwidth links.
@@
Figure 1 depicts the general training stack of the large-scale training cluster in modern corporations. As shown, the operations team manages low-level resources, and the infrastructure team delivers training optimizations[4, 5, 6, 7], with particular emphasis on parallel backbones[8, 9, 10]. Supported by these two teams, multiple algorithm teams focus on adapting LLMs for user-facing applications through various training methods [11, 12]. Notably, since LLMs do not solve all problems, the training cluster also supports other deep learning jobs, such as recommendation models and their specific parallel backbone, TorchRec [13].
Figure 1: The summarized training stack of large-scale training cluster in Ant Group, highlighting XPUTimer’s position.
Efficient distributed LLM training of large scale requires delicate collaboration among the teams across the stack. However, due to the involvement of thousands of software and hardware components, training anomalies frequently arise from various layers. These anomalies include obvious job failures and non-obvious training slowdowns, which should be resolved by different teams.
Specifically, the algorithm teams may unintentionally use incorrect training configurations, leading to training slowdowns. Meanwhile, the slowdowns may also come from unoptimized operators, which fall under the responsibility of the infrastructure team. Moreover, a significant number of job failures are caused by hardware errors. However, the current training stack lacks a dedicated diagnostic framework to address these anomalies. In its absence, resolving such issues requires cross-team investigations, resulting in high communication overhead and low resolution efficiency.
In order to design a deployable diagnostic framework in the training cluster, we identify three main challenges based on real-world operational experiences. C-1: Designing a long-running and lightweight diagnostic framework is challenging. Addressing non-obvious training slowdowns and low-level communication errors requires long-term and in-depth monitoring. However, exhaustive tracing mechanisms, such as PyTorch’s built-in profiler, impose a high runtime memory overhead, making it unsuitable for meeting long-running requirements. C-2: Detecting and diagnosing the underlying root cause is challenging. Errors or slowdowns in LLM training often manifest with similar symptoms, such as process hangs or decreased training speed, which obscure the actual problematic machine or code. C-3: Providing a backbone-extensible diagnostic mechanism is challenging. As shown in Figure 1, at least three backbones are commonly used across the LLM community, not to mention those supporting other deep learning models. Making diagnostic mechanisms extensible to multiple backbones is a significant challenge.
Faced with these challenges, previous efforts[14, 2, 3] targeting distributed LLM training fail to provide comprehensive solutions. This is because these approaches are narrowly designed for specific problems or scenarios, making them inadequate for addressing issues across the software-hardware stack. For example, FALCON [14] and C4D [3] are capable of diagnosing low-level network issues. However, our observations suggest that many slowdowns are caused by upper-level teams, as shown in Table 1. Likewise, tools such as MegaScale [2] are optimized for pre-training scenarios where teams collaborate closely and use only the Megatron backbone, limiting their adaptability to the advancing training stack.
To this end, we present XPUTimer, a real-time and holistic anomaly diagnostic framework designed for efficient distributed LLM training in GPU clusters of thousand-plus scale. XPUTimer consists of two components: a per-training-process tracing daemon and a diagnostic engine. To collect runtime data with low overhead (C-1), the tracing daemon instruments only key code segments rather than blindly tracing all runtime data. Furthermore, as the code segments are carefully selected and intercepted at the level of both Python and C++ runtime, the instrumentation provides sufficient critical information for diagnosis (C-2) while remaining backbone-agnostic (C-3). Notably, XPUTimer is also hardware-extensible, which could provide seamless support for other NPUs.
@@
With the real-time data collected by the tracing daemon, the diagnostic engine detects and diagnoses anomalies, including errors and slowdowns(C-2). For error diagnostics, XPUTimer introduces a novel intra-kernel tracing mechanism, providing fine-grained diagnostics specifically targeting communication-related hang errors. Compared to a binary search using NCCL tests [5, 2], it reduces the complexity of faulty machine diagnostics from $O(\log N)$ to $O(1)$. For slowdown diagnostics, XPUTimer proposes holistic aggregated metrics that encompass not only commonly used macro metrics like training throughput but also novel micro metrics, such as issue latency distribution. With these metrics, XPUTimer can diagnose non-obvious slowdowns (e.g., $2.66\%$) in real-world workloads.
[…]
XPUTimer has been open-sourced at https://github.com/intelligent-machine-learning/dlrover/tree/master/xpu_timer. XPUTimer serves as a core component of DLRover [15], an automated distributed DL system deployed within Ant Group and supported by the LF AI & Data Foundation [16].
Our contributions are as follows.
@@
  • We introduce XPUTimer, a diagnostic framework specifically designed to tackle the critical challenges of long-term monitoring, root cause diagnosis, and backbone extensibility in LLM training diagnostics.
  • We deploy XPUTimer across more than 6,000 GPUs over an 8-month period, deriving typical case studies and practical insights from its daily operations.
Table 1: A comprehensive analysis of common anomalies encountered in Ant Group, annotated with XPUTimer’s primary target.
@@
Network
errors
New
algorithms
Unnecessary
synchronization
Un-optimized
kernels
Memory
management
@@
Slowdown compared to historical
jobs and prior training steps
Slowdown compared to
@@
Team Algorithm Intrastructure Operations Algorithm Intrastructure Operations
Attribution Obvious Obscure Obvious
Comment No need to diagnose XPUTimer’s main target
@@ 2 Background and Motivation
In this section, we introduce the training of LLMs within corporations like Ant Group, and motivate the design of XPUTimer for anomaly diagnostics of LLM training.
@@ 2.1 Large-Scale LLM Training Stack
The pursuit of more powerful large language models [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] has sparked an intense competition between leading corporations. Referring to the software-hardware stack for training LLMs in Figure 1, we delve into how LLMs are reshaping corresponding teams.
@@
With the advancing of LLM algorithms [18], LLMs excel not just in natural language understanding, but also in tackling advanced tasks such as multimodal tasks  [27, 28], reasoning tasks [29, 30]. For instance, customer service teams fine-tune LLMs to develop chatbots that generate accurate responses to complicated questions. Similarly, product development teams leverage multimodal LLMs to generate comprehensive product descriptions by integrating diverse inputs, such as product images and textual data. Consequently, different algorithm teams continue to innovate LLM models [23, 24, 25, 26, 18, 31, 32, 33, 34, 35, 11, 36] tailored for diverse application scenarios.
@@
Training these LLMs is computationally intensive, requiring large-scale GPU clusters of thousand-plus scale [26]. Therefore, leading corporations are continuously investing in large training clusters powered by state-of-the-art hardware [37, 38]. These clusters not only expand in scale but also incorporate the latest GPUs [39, 40] with higher computational performance and advanced interconnect [41, 42]. Efficient operation of these large-scale clusters is critical to the corporations, which requires the operations team to ensure uninterrupted training performance. Its responsibilities include job scheduling, driver updates, hardware maintenance, etc.
@@
To enable easy, efficient, and scalable LLM training within large-scale clusters, the infrastructure teams build the dedicated software stack. This stack bridges the gap between algorithm teams and operations team. Specifically, the infrastructure team focuses on optimizing the software stack by integrating advanced operator libraries [4, 43, 5], training framework [7], and state-of-the-art model parallel backbones [8, 9, 44, 13]. By leveraging this highly optimized training infrastructure, we unlock the full potential of LLMs, enabling innovation and scalability across our various services.
@@ 2.2 Anomalies of Large-scale LLM Training
Due to the complexity of the entire training stack, various issues can easily occur. These issues affect both the training speed of individual training jobs and the overall utilization of training clusters, collectively termed as anomalies. Table 1 presents a distilled analysis of the common anomalies in our real-world cluster, broadly categorized into two primary types: errors and slowdowns. These anomalies often span responsibilities across multiple teams  [2, 3, 14], making them notoriously difficult to diagnose. Thus, resolving these anomalies demands a diagnostic framework that can effectively identify their attributions. However, designing such a framework is far from trivial, posing three key challenges.
Figure 2: Slowdown of a 1024-GPU training job with Llama2-70B under various single-GPU underclocking configurations. Results are evaluated using both Megatron [9] and FSDP [8].
The pain of triggering condition.
Anomalies in large-scale LLM training can arise at any stage and time, as shown in the fourth row of Table 1. First, low-level OS or hardware issues, often attributed to the operations team, may occur randomly, causing runtime hangs or slowdowns. For example, Figure 2 illustrates a significant MFU decline resulting from single-GPU underclocking. Second, training slowdowns caused by codebase modifications may also occur randomly. Specifically, codebase updates on the LLM algorithm typically cause slowdowns at the start, while modifications related to the training infrastructure can result in mid-training slowdowns. A notable example is the slowdown caused by on-demand garbage collection (GC) [2] of Python runtime, as detailed in §5.2.2. In such cases, a long-term background diagnostic mechanism is required to identify these anomalies. However, given the variety of anomalies across the stack, it is challenging to design a lightweight diagnostic framework capable of efficiently collecting key metrics.
The pain of obscured attribution.
[…]
The pain of backbone extensibility.
With decades of advancements in model parallelism for large-scale training, a variety of divergent backbones have been developed to parallelize LLMs. Most LLMs are pre-trained using Megatron [9], a specialized parallel backbone designed for transformer-based LLMs. However, researchers continue to investigate alternative backbones [8, 10, 44] for fine-tuning pre-trained LLMs on downstream tasks or for training more complex large models, such as multimodal models. For instance, fully sharded data parallelism (FSDP)[8] and DeepSpeed[10] are widely utilized within our cluster for training multimodal LLMs. These alternatives are often selected to better align with specific training requirements. However, integrating the diagnostic framework into divergent parallel backbones remains challenging, as their runtime environments and programming interfaces vary significantly.
Existing works have made efforts to build a more robust training stack for large-scale LLM training [2, 3]. They generally operate under several assumptions: (1) training tasks rely on limited backbones like Megatron, (2) all teams work in close collaboration to pre-train a single LLM, or (3) The focus is on specific anomalies, such as communication-related issues. These assumptions fail to reflect reality and these works cannot solve the problems across the full stack. To this end, we propose XPUTimer, a cluster-wide solution deployed in Ant Group’s training cluster. It monitors the real-time progress of LLM training jobs and diagnoses various anomalies precisely.
3 XPUTimer Design
Figure 3: Architecture overview of XPUTimer.
In this paper, we propose XPUTimer, a holistic and real-time diagnostic framework that identifies anomalies in large-scale training with low overhead and pinpoints their corresponding root causes. XPUTimer is designed and implemented following three principles.
  • XPUTimer should be lightweight enough. In this case, it could run as a long-term diagnostic mechanism without compromising the training performance at any scale.
  • XPUTimer should cover enough runtime information. In this case, it could identify the attribution precisely with detailed analysis when encountering anomalies.
  • XPUTimer should be backbone-agnostic. In this case, it is sufficiently extensible to function as a cluster-wide solution for anomaly diagnostics, regardless of the backbones employed.
Architecture overview.

### v2

# Flare: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale

@@ Abstract
The rapid proliferation of large language models has driven the need for efficient GPU training clusters. However, it is challenging due to the frequent occurrence of training anomalies. Since existing diagnostic tools are narrowly tailored to specific issues, there are gaps in their ability to address anomalies spanning the entire training stack. In response, we introduce Flare, a diagnostic framework designed for distributed LLM training at scale. Flare first integrates a lightweight tracing daemon for full-stack and backend-extensible tracing. Additionally, it features a diagnostic engine that automatically diagnoses anomalies, with a focus on performance regressions. The deployment of Flare across 6,000 GPUs has demonstrated significant improvements in pinpointing deficiencies in real-world scenarios, with continuous operation for over eight months.
@@ 1 Introduction
The advent of large language models (LLMs) has revolutionized the deep learning training community, driving substantial advancements in artificial intelligence-generated content (AIGC). Recognizing their transformative potential to enhance user experiences, leading corporations are proactively leveraging LLMs to enhance a wide array of user-oriented services [antgroup, jiangMegaScaleScaling, dongEnhancingLargescale]. To meet the computational demands of LLM training, they construct large-scale training clusters comprising the latest GPUs interconnected via high-bandwidth links.
@@
Figure 1 depicts the general training stack of the large-scale training cluster in modern corporations. As shown, the operations team manages low-level resources, and the infrastructure team delivers training optimizations[cublas2024, nccl2024, daoFlashAttentionFast, paszkePyTorchImperative], with particular emphasis on parallel backends[zhaoPyTorchFSDP, shoeybiMegatronLMTraining, rajbhandariZeROMemory]. Supported by these two teams, various algorithm teams focus on adapting LLMs for user-facing applications through various training methods [shengHybridFlowFlexible, hanParameterEfficientFineTuning]. Notably, the training cluster also supports other deep learning jobs, such as recommendation models and their specific parallel backend, TorchRec [ivchenkoTorchRecPyTorch].
Figure 1: The training stack of large-scale training cluster.
Efficient large-scale distributed LLM training requires effective coordination of software and hardware across the stack. However, training anomalies frequently arise from various layers. These anomalies include obvious job failures and non-obvious training slowdowns, which can be categorized into three types: performance regressions, job failures, and fail-slows. Performance regressions may result from incorrect training settings (addressed by algorithm teams) or unoptimized operators (addressed by infrastructure teams). In contrast, most job failures and fail-slows arise from hardware issues, which are handled by the operations team. One key insight in this paper is that fail-slows are sudden and easy-to-detect slowdowns due to transient issues in components, whereas performance regressions are persistent and hard-to-detect slowdowns caused by code updates or configuration drift.
Drawing on our daily operational experiences, teams have spent significant time in cross-team collaborations to address these anomalies, often repeating efforts on similar cases. For instance, performance regressions caused by unnecessary synchronizations introduced by algorithm teams often require collaboration with the infrastructure team. Yet, eliminating these synchronizations usually only involves deleting a few lines of code in the training scripts. Therefore, a dedicated diagnostic framework is urgently needed to mitigate this.
The framework first should collect sufficient runtime data for anomaly detection and potential causes identification, enabling teams to resolve them with minimal collaboration. To this end, runtime data must be continuously gathered from all layers of the training stack, especially from the software in the algorithm and infrastructure teams. Secondly, root causes should be narrowed as much as possible; otherwise, even if recurrent anomalies are routed to the right teams, they may still be unable to resolve them independently.
To design such a diagnostic framework, we identify two main challenges. C-1: Designing a full-stack tracing mechanism with backend extensibility is challenging. The tracing mechanism collects runtime data for anomaly diagnosis. While full-stack tracing is needed to capture sufficient runtime data, diverse parallel backends calls for backend-extensibility, However, these requirements often conflict, as full-stack tracing usually intrudes into parallel backend codebases, reducing backend-extensibility. C-2: Detecting and diagnosing root causes is challenging. Errors or fail-slows in LLM training often show similar symptoms, such as process hangs or sudden drops in training speed. Worse, regressions cannot be directly confirmed by comparing macro metrics, such as current training throughput, with historical runs, since throughput declines may be expected. These factors make it difficult to detect anomalies and narrow down their root causes.
Faced with these challenges, previous efforts[wuGREYHOUNDHunting, jiangMegaScaleScaling, dongEnhancingLargescale] targeting distributed LLM training fail to provide comprehensive solutions. This is because they are narrowly designed for specific problems or scenarios, making them inadequate for addressing issues across the software-hardware stack. For example, Greyhound [wuGREYHOUNDHunting], C4D [dongEnhancingLargescale], and Holmes [yaoHolmesLocalizing] can only trace and diagnose errors or fail-slows under the responsibility of operations team. However, they could not solve the performance regressions originated from upper-level teams. Besides, tools such as MegaScale [jiangMegaScaleScaling] are optimized for pre-training scenarios involving a single algorithm team and a single backend. It intrudes into the backend codebase to achieve full-stack tracing, but this tight coupling makes it difficult to plug into other parallel backends. Moreover, MegaScale only provides visualization of distributed training and relies on cross-team collaboration and manual effort to pinpoint root causes for detected anomalies.
To this end, we present Flare, an anomaly diagnostic framework for divergent LLM training in GPU clusters at thousand-plus scale. It places particular emphasis on diagnosing performance regression anomalies from algorithm and infrastructure teams. Flare consists of two components: a per-training-process tracing daemon and a diagnostic engine. To collect sufficient runtime data with backend-extensibility (C-1), the tracing daemon selectively instruments key code segments in a plug-and-play manner across both Python and C++ runtimes. Most importantly, it leverages CPython tracing mechanisms to trace Python code segments without intruding into backbone codebases. It also provides an easy-to-play interface—simply adding shell environments—for extending to new backends and allowing all teams to selectively trace Python code segments.
@@
With real-time data collected by the tracing daemon, the diagnostic engine detects and diagnoses anomalies while narrowing their root causes (C-2). As for error diagnostics, Flare introduces a novel intra-kernel inspecting mechanism, providing fine-grained diagnostics specifically targeting communication-related hang errors. It enables $\mathcal{O}(1)$-complexity faulty machine diagnostics without exhaustive or blind searches based on NCCL tests. As for slowdown diagnostics, Flare proposes holistic aggregated metrics that encompass not only commonly used macro metrics like training throughput but also novel micro metrics, such as issue latency distribution. Importantly, the new proposed micro metrics enable Flare to detect and diagnose regressions. Overall, with these metrics, Flare enables teams to independently address those routed slowdowns, including subtle ones such as a $2.66\%$ regressions in real-world workloads.
We extensively evaluated Flare in terms of its runtime overhead. Flare incurs an average latency overhead of only $0.43\%$ across various LLMs and backends on 1024 H800 GPUs. Meantime, Flare only generates just $1.5\text{MB}$ of tracing logs per GPU in a real-world model trained on 1536 H800 GPUs. Flare has been deployed in our training cluster at Ant Group[antgroup], utilizing over 6,000 GPUs to date. The greatest benefit of Flare’s deployment is that it eases the handling of recurrent anomalies. Flare has been open-sourced at DLRover. It serves as a core component of DLRover [dlrover], an automated distributed DL system deployed within Ant Group and supported by the LF AI & Data Foundation [lfaidata]. Our contributions are as follows.
@@
  • We introduce Flare, a diagnostic framework specifically designed to tackle the critical challenges of full-stack tracing, root cause diagnosis, and backend extensibility in LLM training diagnostics.
  • We deploy Flare across more than 6,000 GPUs over an 8-month period, deriving typical case studies and practical insights from its daily operations.
Table 1: A comprehensive analysis of anomalies encountered in Ant Group, with Flare’s primary target highlighted in yellow.
@@
[…]
@@
[…]
Fail-slows compared to
@@
[…]
@@ 2.1 Large-Scale LLM Training Stack
Referring to the software-hardware stack in Figure 1, we delve into how LLMs are reshaping teams in leading corporations.
@@
With the advancing of LLM algorithms [deepseekmoe], LLMs excel not just in natural language understanding, but also in tackling advanced tasks such as multimodal tasks  [vit, sora], reasoning tasks [openaio1, rstar]. For instance, customer service teams fine-tune LLMs to develop chatbots that generate accurate responses to complicated questions. Similarly, product development teams leverage multimodal LLMs to generate comprehensive product descriptions by integrating inputs, such as product images and textual data. Consequently, different algorithm teams continue to innovate LLM models tailored for diverse application scenarios.
@@
Training these LLMs is computationally intensive, requiring large-scale GPU clusters of thousand-plus scale [llama3-1]. Therefore, leading corporations are continuously investing in large training clusters powered by state-of-the-art hardware [xai_colossus_training_cluster, meta_ai_training_infrastructure]. These clusters not only expand in scale but also incorporate the latest GPUs [nvidia_ampere_architecture, nvidia_hopper_architecture] with higher computational performance and advanced interconnect [nvidia_nvswitch, nvidia_nvl72]. Efficient operation of these large-scale clusters is critical to the corporations, which requires the operations team to ensure uninterrupted training performance.
@@
To enable easy, efficient, and scalable LLM training within large-scale clusters, the infrastructure teams build the dedicated software stack. This stack bridges the gap between algorithm teams and operations team. Specifically, the infrastructure team focuses on optimizing the software stack by integrating advanced operator libraries [cublas2024, cutlass2024, nccl2024], training framework [paszkePyTorchImperative], and state-of-the-art model parallel backends [zhaoPyTorchFSDP, shoeybiMegatronLMTraining, liangTorchTitanOnestop, ivchenkoTorchRecPyTorch]. Parallel backends are fundamental for efficiently training LLMs, enabling tensor, data, and pipeline parallelism across thousands of GPUs.
@@ 2.2 Anomalies of Large-scale LLM Training
Due to the complexity of the entire training stack, various issues can easily occur. These issues affect both the training speed of individual training jobs and the overall utilization of training clusters, collectively termed as anomalies.
Table 1 presents a distilled analysis of the common anomalies in our real-world cluster, broadly categorized into two primary types: runtime errors and slowdowns. Based on symptom differences, slowdowns can be further divided into two types: performance regressions (persistent slowdowns) and fail-slows (sudden and acute slowdowns).
Based on three months of traces from a cluster with 6000+ GPUs, there are 127 errors and 135 slowdowns observed across 3047 jobs. The slowdowns include 78 performance regressions and 57 fail-slows. When such anomalies occur, they are first reported by algorithm teams, and then multiple teams coordinate and spend significant effort to resolve them.
Specifically, for runtime errors and fail-slows, the operations team intervenes to resolve these acute and easy-to-detect job failures. Without timely intervention, training jobs may fail or run at very low throughput, significantly reducing training efficiency. For performance regressions, algorithm teams are often unfamiliar with why these regressions occur with their own code changes, so collaboration with infrastructure teams is required to resolve them. In a training stack like Figure 1, the situation worsens: the infrastructure team supports multiple algorithm teams, each potentially using different parallel backends for LLM training. As recurrent regressions arise across different algorithm teams, the infrastructure team must repeatedly expend effort to address them.
This dilemma motivates a dedicated diagnostic framework to streamline anomaly resolution, especially performance regressions. To achieve this, we must collect sufficient runtime data across the training stack. For example, overall training throughput helps detect fail-slows, while per-operator performance reveals regressions introduced by minor software changes. We must also seamlessly support multiple parallel backends used by different algorithm teams. Finally, for detected anomalies, we aim to narrow down their root causes as much as possible, enabling them to be routed to the appropriate team and resolved independently without unnecessary cross-team collaboration.
However, designing such a framework is far from trivial, posing two key challenges.
The difficulty of full-stack, backend-extensible tracing.
Diagnosing performance regressions across diverse parallel backends requires full-stack and backend-extensible tracing to collect sufficient runtime data. These goals often conflict. Full-stack tracing tends to be intrusive to each backend’s codebase, whereas extensibility demands minimal changes to the backends. A typical example is MegaScale [jiangMegaScaleScaling], which enables low-overhead tracing for FSDP–the native PyTorch backend–by patching the FSDP codebase. Extending support to other backends would require similar patches. A tracing design for runtime data collection in training jobs of multiple algorithm teams should avoid such backend-intrusive modifications.
The difficulty of narrowing down attribution.
Errors and fail-slows in Table 1 often show obvious symptoms, as the job either fail or show sudden training speed degradation. There are various works like C4D [dongEnhancingLargescale], Greyhound [wuGREYHOUNDHunting], and Holmes [yaoHolmesLocalizing] that are proposed to diagnose them. However, automatic performance regressions are rarely explored in prior work. Attributing them is more challenging than diagnosing fail-slows. First, they are harder to detect. Regressions are persistent slowdowns caused by software changes, and monitoring training throughput alone is insufficient. Algorithm teams may not notice them and may assume jobs are running normally. Second, linking a detected regression to the corresponding software change is difficult. For example, whether Python garbage collection causes a regression depends on how it is invoked. Thus, new metrics must be designed based on operational expertise, especially from infrastructure team, to detect and diagnose regressions.
3 Flare Design
Figure 2: Architecture overview of Flare.